# Improving Sensitivity of Vector-like Top Partner Searches with Jet Substructure

### v1

@@ IV Event reconstruction and selection
Events having a minimum of one reconstructed pp interaction vertex are retained for this study. The vertices must be reconstructed from at least four tracks that have a longitudinal distance $|\text{d}_{\text{z}}|<24$ cm and a radial distance $|\text{d}_{\text{xy}}|<2$ cm from the nominal interaction point. If multiple vertices are found in an event, the one with the highest value of summed $\text{p}_{\text{T}}^{2}$ of clustered particles is taken as the primary pp interaction vertex. In contrast, the others are identified as the pileup vertices. Individual particles are reconstructed via the particle-flow algorithm Sirunyan et al. (2017). Muons (Electrons) are required to have $\text{p}_{\text{T}}\geqslant 30\ (35)$ GeV, $|\eta|\leqslant 2.4\ (2.1)$ and to pass the “tight” identification criteria Sirunyan et al. (2018b, 2021). The selected muon and electron candidates must pass the criterion on relative isolation ($\text{I}_{\text{rel}}$) to be within $6\%$ of their respective transverse momentum. Additionally, the selected electron candidate is required to satisfy $|\text{d}_{\text{xy}}|\leqslant 0.05\ (0.1)$ cm and $|\text{d}_{\text{z}}|\leqslant 0.1\ (0.2)$ cm relative to the primary vertex in the barrel (endcap) region of the electromagnetic calorimeter (ECAL). Events containing additional muons (electrons) with $\text{p}_{\text{T}}\geqslant 10\ (15)$ GeV, $|\eta|\leqslant 2.4\ (2.5)$, and passing the “loose” (“veto”) identification criteria are rejected. In such cases, the criteria on the additional lepton isolation are relaxed to $\text{I}_{\text{rel}}\leqslant 20\%$ for muons and $\text{I}_{\text{rel}}\leqslant 18\%\ (16\%)$ for electrons in the barrel (endcap) ECAL region. Correction factors for lepton reconstruction efficiencies are applied to simulated events for both selected and vetoed leptons in bins of lepton $\text{p}_{\text{T}}$ and $\eta$; to mitigate observed differences with collision data Sirunyan et al. (2018b, 2021). Standard jets are reconstructed using the AK clustering algorithm Cacciari et al. (2008) with radii of 0.4 and 0.8 for the narrow (AK4) and fat (AK8) jets respectively, as implemented in the FastJet package Cacciari et al. (2012). The effect of additional tracks and energy deposits in calorimeters from pileup on the AK4 and AK8 jet momenta are mitigated by discarding tracks identified to be originating from pileup vertices, as well as by applying an offset correction to account for residual neutral pileup contributions Cacciari and Salam (2008); Sirunyan et al. (2020). “Tight lepton veto” identification criteria CMS (2017) are applied to suppress jets arising from spurious sources, such as electronic noise in the calorimeters and jets containing “tight”-ly identified muons or electrons. The combined secondary vertex tagging algorithm version 2 (CSVv2Sirunyan et al. (2018c) is used to identify AK4 jets originating from bottom quarks or antiquarks. Candidate events are required to have exactly one AK8 jet with $\text{p}_{\text{T}}$ $\geqslant 200~{}$GeV and within $|\eta|\leqslant 4.0$, as mentioned in Section II.2. In addition to the selected AK8 jet, events must have at least one non-overlapping AK4 jet with $\text{p}_{\text{T}}\geqslant 30$ GeV, $|\eta|\leqslant 2.5$ and passing the “medium” working point defined on the CSVv2 score. Finally, events should pass $\text{p}_{\text{T}}^{\text{miss}}\geqslant 30~{}$GeV criterion.
In addition to the AK4 and AK8 jets, we reconstruct the anti-$\text{k}_{\text{T}}$ DR jets (see Section III.1) for the selected events [^1] The jet selection criteria are applied to AK4 and AK8 jets, and not to dynamic radius jets. This choice is made primarily due to the absence of optimized pileup contamination removal, jet identification working points, jet energy resolution and correction etc. for the dynamic radius jets. using the particle-flow candidate information as input. The DR algorithm is used to form two jet collections per event, corresponding to $R_{0}$ values of $0.3$ and $0.7$ for narrow and fat jets, respectively. The narrow jets that do not overlap with any fat jet in the event, are considered as independent. The CSVv2 score of the nearest AK4 jet is assigned as the b-tagging discriminant value for each narrow jet.
Table 5 lists signal and background yields after all event selection steps for the two event categories. Figures 6 and 7 show several kinematic distributions after applying all selection criteria in $e~{}$+ jets and $\mu~{}$+ jets final states, respectively.
@@
$\mathrm{t\bar{t}}$ + t($\bar{\text{t}}$)W + $\mathrm{t\bar{t}Z}$ $\mathrm{t\bar{t}W}$ WV + jets QCD
jets jets (V=~{}=~{}=W or Z) multijet
$e~{}$+ jets 18313 29544 620 825 4077 35634 89013 25
@@
Figure 6: Distributions of various kinematic and event shape observables in the $e$ + jets final state.
Figure 7: Distributions of various kinematic and event shape observables in the $\mu$ + jets final state.
@@ V Results
Several observables are combined into a boosted decision tree (BDT) discriminant per event category to separate signal from the collective SM backgrounds. These are listed in Appendix A along with the respective ranks based on their importance. The simulated events are separated into two independent sub-samples, one for training and the other for validation, in $70\%:30\%$ ratio. An event is picked randomly to belong to any one of these sub-samples. Each process is weighted according to its purity during training. Two independent trainings are performed, one for fixed-radius jets and another for DR jets. The adaptive boosting algorithm Freund and Schapire (1997) implemented within the TMVA package Hoecker et al. (2007) has been used. A check for overtraining is performed by comparing the respective BDT scores of signal and background between training and validation sub-samples. The corresponding Kolmogorov-Smirnov probabilities range between $32\%$$97\%$, indicating no overtraining. The performance of the BDTs is studied via the receiver-operator characteristic (ROC) curves for different thresholds applied to the fat jet $\text{p}_{\text{T}}$. Figure 8 shows the BDT scores of signal and background as well as comparisons of ROC curves for different fat jet $\text{p}_{\text{T}}$ thresholds for fixed-radius and DR jet algorithms. The BDT performance falls off relatively slower for tighter thresholds on fat jet $\text{p}_{\text{T}}$ in the case of DR jet algorithm as indicated by the ROC curves.
Figure 8: BDT scores (upper row) along with the ROCs (lower row) for fixed-radius (left) and dynamic radius (right) jets.
@@
Selection thresholds on the BDT scores are optimized based on the so-called “Punzi figure-of-merit” ($\text{FOM}_{\text{Punzi}}$) defined in Ref. Punzi (2003). The variation of the signal significance defined as $\frac{S}{\sqrt{S+B}}$ and the $\text{FOM}_{\text{Punzi}}$ are studied with fat jet $\text{p}_{\text{T}}$ threshold for the fixed-radius and DR jet algorithms after applying a selection on the BDT scores $\geqslant 0.3$. From Fig. 9, it is evident that the DR jet algorithm outperforms the fixed-radius one in the highly boosted (fat jet $\text{p}_{\text{T}}\geqslant 750~{}$GeV) region of the phase space. The adaptive approach adopted by the DR algorithm (see Section III.1) ensures that all daughters of the $T$ quark are essentially captured within the fat jet by optimizing the jet size, which in turn leads to improved reconstruction of the signal events, in this region. The DR algorithm shows higher sensitivity towards large $M_{T}$ values (see Fig. 3) compared to the fixed-radius algorithm, allowing for more phase space to hunt for the signal process.
Figure 9: Comparison of $\text{FOM}_{\text{Punzi}}$ (left) and signal significance (right) between fixed-radius and dynamic radius jets with the fat jet $\text{p}_{\text{T}}$ threshold.
@@
We then calculate expected $2\sigma$ upper limits on the chromomagnetic couplings, assuming $C_{tRL}=C_{tLR}$, for the top partner masses in the range 1.4 TeV to 3 TeV. For each mass point, the upper limit has been obtained by varying the coupling strength such that the signal significance reaches $2\sigma$. Figure 10 shows the resulting upper limits for the three different fat jet $\text{p}_{\text{T}}$ thresholds. In the plots, the shaded regions correspond to values of the coupling excluded by our analysis at $95\%$ CL under the assumption of no signal observation. The upper limits are estimated using both fixed-radius and DR jet algorithms. In all cases, the expected limit from the DR algorithm is smaller than that from the fixed radius one, indicating that the analysis with dynamic radius jets has better sensitivity to smaller coupling values. The improvement is more prominent for the fat jet $\text{p}_{\text{T}}$ threshold of 900 GeV, suggesting that the dynamic radius jet algorithm is especially effective in the high fat jet $\text{p}_{\text{T}}$ regime.
@@
Figure 10: Expected $2\sigma$ upper limit on the chromomagnetic coupling $C_{tRL}=C_{tLR}$ as a function of the top partner mass. The three panels correspond to three different fat jet $\text{p}_{\text{T}}$ thresholds. For these plots, we use $\Lambda=8$ TeV, and all other parameters are kept the same as in the benchmark point defined in Eq. (7). Shaded regions indicate the parameter space excluded by our analysis, assuming no signal is observed.
@@ VI Summary and Outlook
In this work, we performed a thorough investigation of collider prospects of a vectorlike top partner produced in association with a standard model top quark via a chromomagnetic coupling. The top partner is assumed to decay into the following channels: $tZ$, $th$, $tg$, or $bW$. To effectively suppress the huge QCD background, we focused on semileptonic final states, i.e., events with exactly one isolated lepton arising either from the top quark or from the heavy boson decay. Depending on the decay modes of the top partner, the final state could contain one or two fat jets along with a single lepton and $b$-tagged jet(s). We performed the analysis in the event category that contains single fat jet ($\ell$ + 1 fat jet + $\geqslant 1$ $b$-tagged jet) due to its comparatively better signal-to-background ratio.
We carried out a detailed collider analysis at the 13 TeV LHC with an integrated luminosity of 300 fb-1, incorporating realistic pileup conditions. For signal events, the Madgraph5_aMC@NLO + Pythia8 + Delphes simulation chain has been employed; while for background estimation, appropriate simulated samples from CMS Open Data are utilized. To enhance signal discrimination, we utilized state-of-the-art jet substructure tools such as Soft Drop grooming, $N$-subjettiness, and Energy Correlation Functions, along with several event shape observables like Fox-Wolfram moments, Sphericity, and Aplanarity. Multivariate discriminants based on boosted decision trees are developed for better signal-to-background separation. We then optimized a selection thresholds on these discriminants based on the Punzi figure-of-merit to improve signal sensitivity.
The result is presented as the upper limit on the chromomagnetic coupling between the top quark and its heavy vectorlike partner. Assuming the cut off scale of $\Lambda=8$ TeV, the upper limit on the coupling is found to lie within 0.2 to 1.5 for the top partner mass range of 1.4 to 3 TeV.

### v2

@@ IV Event Reconstruction and Selection
Various key features of the event reconstruction and selection are presented below.
Selection Criteria
@@
Events having a minimum of one reconstructed pp interaction vertex are retained for this study. The vertices must be reconstructed from at least four tracks that have a longitudinal distance $|\text{d}_{\text{z}}|<24$ cm and a radial distance $|\text{d}_{\text{xy}}|<2$ cm from the nominal interaction point. If multiple vertices are found in an event, the one with the highest value of summed $\text{p}_{\text{T}}^{2}$ of clustered particles is taken as the primary pp interaction vertex. In contrast, the others are identified as the pile up vertices. Individual particles are reconstructed via the particle-flow algorithm CMS:2017yfk . Muons (Electrons) are required to have $\text{p}_{\text{T}}\geqslant 30\ (35)$ GeV, $|\eta|\leqslant 2.4\ (2.1)$ and to pass the “tight” identification criteria CMS:2018rym ; CMS:2020uim . The selected muon and electron candidates must pass the criterion on relative isolation ($\text{I}_{\text{rel}}$) to be within $6\%$ of their respective transverse momentum. Additionally, the selected electron candidate is required to satisfy $|\text{d}_{\text{xy}}|\leqslant 0.05\ (0.1)$ cm and $|\text{d}_{\text{z}}|\leqslant 0.1\ (0.2)$ cm relative to the primary vertex in the barrel (endcap) region of the electromagnetic calorimeter (ECAL). Events containing additional muons (electrons) with $\text{p}_{\text{T}}\geqslant 10\ (15)$ GeV, $|\eta|\leqslant 2.4\ (2.5)$, and passing the “loose” (“veto”) identification criteria are rejected. In such cases, the criteria on the additional lepton isolation are relaxed to $\text{I}_{\text{rel}}\leqslant 20\%$ for muons and $\text{I}_{\text{rel}}\leqslant 18\%\ (16\%)$ for electrons in the barrel (endcap) ECAL region. Standard jets are reconstructed using the AK clustering algorithm Cacciari:2008gp with radii of 0.4 and 0.8 for the narrow (AK4) and fat (AK8) jets respectively, as implemented in the FastJet package Cacciari:2011ma with the particle-flow candidates as inputs. The effect of additional tracks and energy deposits in calorimeters from pile up on the AK4 and AK8 jet momenta are mitigated by discarding tracks identified to be originating from pile up vertices, as well as by applying an offset correction to account for residual neutral pile up contributions Cacciari:2007fd ; CMS:2020ebo . “Tight lepton veto” identification criteria CMS-PAS-JME-16-003 are applied to suppress jets arising from spurious sources, such as electronic noise in the calorimeters and jets containing “tight”-ly identified muons or electrons. The combined secondary vertex tagging algorithm version 2 (CSVv2CMS:2017wtu is used to identify AK4 jets originating from bottom quarks or antiquarks. Candidate events are required to have exactly one AK8 jet with $\text{p}_{\text{T}}$ $\geqslant 200$~GeV and within $|\eta|\leqslant 4.0$, as mentioned in Section II.2. In addition to the selected AK8 jet, events must have at least one non-overlapping AK4 jet with $\text{p}_{\text{T}}\geqslant 30$ GeV, $|\eta|\leqslant 2.5$ and passing the “medium” working point defined on the CSVv2 score. Finally, events should pass $\text{p}_{\text{T}}^{\text{miss}}\geqslant 30$~GeV criterion.
We note that a full detector simulation with Geant4 with CMS detector specification has been used in the background MC samples. However, for the signal MC samples, a fast detector simulation using Delphes has been used. According to Ref. deFavereau:2013fsa , the overall agreement in the average performance of Delphes with the full simulation of the CMS detector based on Geant4 is reasonably good for different high-level physics objects, especially in the phase space region of interest. Furthermore, the measured efficiencies of lepton (both selected and vetoed leptons) CMS:2018rym ; CMS:2020uim and jet CMS-PAS-JME-16-003 ; CMS:2017wtu reconstruction are incorporated into the Delphes for signal events in bins of their respective $\text{p}_{\text{T}}$ and $\eta$. In case of backgrounds, to mitigate observed differences with collision data; respective efficiency corrections are applied CMS:2018rym ; CMS:2020uim ; CMS-PAS-JME-16-003 ; CMS:2017wtu . Table 6 lists signal and background yields after all event selection steps for the two lepton flavours, while Figures 68 show the distributions of observables related to lepton, jet, and event kinematics as indicated in Table 8.
@@
$\mathrm{t\bar{t}}$ tW $\bar{\text{t}}$W $\mathrm{t\bar{t}Z}$ $\mathrm{t\bar{t}W}$ WV + jets QCD
+ jets + jets (V=~=~W or Z) multijet
ee~+ jets 18313 29544 620 825 4077 35634 89013 25
@@
Reconstruction of Dynamic Radius Jets
In addition to the AK4 and AK8 jets, we reconstruct the DR anti-$\text{k}_{\text{T}}$ jets (see Section III.1) for the selected events [^4] The jet selection criteria are applied to AK4 and AK8 jets, and not to dynamic radius jets. This choice is made primarily due to the absence of optimized pile up contamination removal, jet identification working points, jet energy resolution and correction etc. for the dynamic radius jets. using the particle-flow candidate information as input. The DR algorithm is used to form two jet collections per event, corresponding to $\rm R_{0}$} values of $0.3$ and $0.7$ for narrow and fat jets, respectively. The narrow jets that do not overlap with any fat jet in the event, are considered for further analysis. The CSVv2 score of the nearest AK4 jet is assigned as the b-tagging discriminant value for each narrow jet.
Signal Event Reconstruction with Dynamic Radius Jets
It is important to understand the impact of DR clustering method on the signal event reconstruction, especially in the boosted regime. A simple study is performed for this purpose by applying the following selections on the jet substructure observables shown in Figure 5, to identify jets due to hadronically decaying boosted W or Z bosons and top quarks in the simulated signal events.
[…]
The results are presented in Table 7 in terms of the selection efficiencies of the boosted W or Z bosons and top quarks for the two clustering methods at high fat jet $\text{p}_{\text{T}}$. Higher selection efficiencies are observed for both the boosted W- or Z-like and top-like jets in the case of DR clustering method, indicating better reconstruction of the daughters of the $T$ quark in the boosted regime; thus improving the overall signal event reconstruction in this region. The adaptive approach adopted by the DR clustering (see Section III.1) helps in capturing all daughters of the boosted W or Z boson or the top quark originating from the $T$ quark decay (see Table 2) within the fat jet by optimizing the jet size.
Table 7: Selection efficiencies of boosted W- or Z-like and top-like jets for fat jet $\text{p}_{\text{T}}$ $\geqslant 900$~GeV corresponding to fixed- and dynamic radius clustering methods in simulated signal events.
Process Clustering Selection efficiency
method Boosted W or Z boson Boosted top quark
Signal Fixed-radius $38.3\%$ $46.5\%$
Dynamic radius 41.1% 47.5%
Figure 6: Distributions of lepton $\text{p}_{\text{T}}$ (upper), fat jet $\text{p}_{\text{T}}$ (middle), leading narrow jet $\text{p}_{\text{T}}$ (lower), in the $e$ + jets (left) and $\mu$ + jets (right) final states.
[…]
Figure 8: Distributions of event shape observables, namely, $10^{\text{th}}$ order Fox–Wolfram moment (upper), Sphericity (middle), and Aplanarity (lower), in the $e$ + jets (left) and $\mu$ + jets (right) final states.
@@ V Results
V.1 Multivariate Analysis
Several observables are combined into boosted decision tree (BDT) discriminants to effectively separate signal from the collective SM backgrounds listed in Table 3. The input observables to the BDT discriminants are listed in Table 8. The most important aspects of the BDT analysis are discussed in details below.
Table 8: List and description of input observables to the BDT discriminants.
Type Observable Description
[…]
$|\eta_{\ell}|$ Lepton $|\eta|$
[…]
$|\eta_{\text{j}}|$ $|\eta|$ of the leading narrow jet that is not part of any fat jet
[…]
$|\eta_{\text{fj}}|$ Leading fat jet $|\eta|$
[…]
$\tau_{21}$ 2-to-1 subjettiness ratio of the leading fat jet
$\tau_{32}$ 3-to-2 subjettiness ratio of the leading fat jet
Event kinematics $\text{p}_{\text{T}}^{\text{miss}}$ Missing transverse momentum
[…]
$\text{H}_{\text{T}}$ Scalar $\text{p}_{\text{T}}$ sum of non-overlapping narrow and fat jets
$_{\text{jets}} N$umber of narrow jets that are not part of the fat jet
$_{\text{b-jets}} N$o. of narrow b-jets that are not part of any fat jet
[…]
S Sphericity obtained from the eigenvalues of the momentum tensor ATLAS:2012tch
A Aplanarity obtained from the eigenvalues of the momentum tensor ATLAS:2012tch
Choice of Inputs
A variety of jet substructure and event observables have been studied for this purpose, some of which are shown in Figures 48. However, a subset of these are chosen as inputs to the BDT discriminants since there are large correlations between some of the pairs. Observables that were highly correlated with each other were removed from the list of inputs to keep only the less correlated ones. For example, none of the ECFs or their ratios have been used as the inputs to the BDT discriminants because they would be highly correlated to $N$-subjettiness and their ratios. Different observables have different capabilities of discriminating the signal process from the SM backgrounds which can be broadly classified into two types, namely reducible and irreducible. The backgrounds having top quarks in the final state compose the irreducible type; while the others constitute the reducible type. Boosted decision trees are fully tree-based, human-readable structure with a finite maximum depth that selects per tree only a few from the input observables list which provide the best discrimination between signal and backgrounds. Thus, using a relatively large number of observables as inputs to the BDTs does not introduce any bias; instead it mostly increases the computational cost due to the residual correlations among the observables, even in the worst case where the performance is not significantly improved by including an observable. This strategy has been widely adopted in experimental searches such as Ref. CMS:2022fck as well as in phenomenological studies such as Ref. Ghosh:2025gue . The input observables are ranked according to their so-called “importance” to the BDT discriminants as shown in Fig. 12 in Appendix B. The “importance” of an observable is derived by counting how often that observable has been used to split decision tree nodes, and by weighting each split occurrence by the separation gain-squared it has achieved and by the number of events in the node Breiman:2017lcz .
Training Setup
@@ A Aplanarity obtained from the eigenvalues of the momentum tensor ATLAS:2012tch
The simulated events are separated into two independent sub-samples, one for training and the other for validation, in $70\%:30\%$ ratio. An event is picked randomly to belong to any one of these sub-samples. The adaptive boosting algorithm Freund:1997xna implemented within the TMVA package Hocker:2007ht has been used to train the discriminants keeping the learning rate, minimum node size, and maximum depth set to $0.5$, $2.5\%$, and $3$, respectively. Two independent trainings are performed, one for fixed-radius jets and another for DR jets, with exactly same settings. In each training, samples are selected after applying a $\text{p}_{\text{T}}$ threshold of $200$ GeV on the leading fat jet as indicated in Section IV. Each process is weighted according to its purity, defined as $\frac{N_{i}}{\sum_{i}N_{i}}$ with $i$ running over the signal and background processes, during training. Checks for overtraining are performed by comparing the respective BDT scores of signal and background between training and validation sub-samples. The corresponding Kolmogorov-Smirnov probabilities Chakravarti_Laha_Roy_1967 range between $32\%$$97\%$, indicating no overtraining. In the left column of Fig. 9, we show the BDT score distributions of signal and background corresponding to fixed-radius (upper left) and DR (lower left) algorithms for the nominal event selection.
Performance Evaluation
We evaluate the BDT scores also for events passing different thresholds on the leading fat jet $\text{p}_{\text{T}}$, and the performances of the BDT discriminants are then studied via the receiver-operator characteristic (ROC) curves for each of these thresholds separately, along with the nominal case. This exercise is primarily aimed at understanding the sensitivity in the high fat jet $\text{p}_{\text{T}}$ regime. The ROC curves corresponding to different fat jet $\text{p}_{\text{T}}$ thresholds are shown in the right column of Fig. 9 for fixed-radius (upper right) and DR (lower right) clustering methods. The BDT performance falls off less rapidly for tighter thresholds on fat jet $\text{p}_{\text{T}}$ in the case of DR jet clustering as indicated by the ROC curves in Fig. 9.
Figure 9: BDT scores (left) along with the ROC curves (right) for fixed-radius (upper) and dynamic radius (lower) jets. The BDT score distributions are obtained for the nominal fat jet $\text{p}_{\text{T}}$ threshold of $200$ GeV (see Section IV) which corresponds to the green curves on the right column on each row. The other ROC curves are obtained by applying an additional criteria on the fat jet $\text{p}_{\text{T}}$ as indicated in the legend.
V.2 Signal Sensitivity and Interpretation
The BDT discriminants are then utilized for a comparison of signal sensitivities followed by interpretations in terms of expected limits on the plane spanned by $T$ quark mass and its chromomagnetic coupling, corresponding to the two clustering approaches.
Comparative Study of Signal Sensitivities
@@ V.2 Signal Sensitivity and Interpretation
Selection thresholds on the BDT scores are optimized based on the so-called “Punzi figure-of-merit” ($\text{FOM}_{\text{Punzi}}$) defined in Ref. Punzi:2003bu . The signal sensitivity is characterised via two different metrics, namely the signal significance defined as $\frac{S}{\sqrt{S+B}}$ and the $\text{FOM}_{\text{Punzi}}$, to check the consistency and robustness of the conclusion. The variation of these metrics are studied with the fat jet $\text{p}_{\text{T}}$ threshold for the fixed-radius and DR jet clustering methods after applying the optimized selection on the BDT scores $\geqslant 0.3$. From Figure 10, it is evident that the DR jet algorithm outperforms the fixed-radius one in the highly boosted (fat jet $\text{p}_{\text{T}}\geqslant 750$~GeV) region of the phase space. The DR algorithm shows higher sensitivity towards large $M_{T}$ values (see Fig. 3) compared to the fixed-radius algorithm, allowing for more phase space to hunt for the signal process.
Interpretation via Expected Limits
@@
The result of the signal sensitivity study is interpreted by calculating the expected $2\sigma$ upper limits on the chromomagnetic couplings, assuming $C_{tRL}=C_{tLR}$, for the top partner masses in the range 1.4 TeV to 3 TeV. For each mass point, the expected upper limit has been obtained by varying the coupling strength such that the signal significance reaches $2\sigma$. Figure 11 shows the resulting upper limits for the three different fat jet $\text{p}_{\text{T}}$ thresholds. In these plots, the shaded regions correspond to values of the coupling excluded by our analysis at $95\%$ CL under the assumption of no signal observation. The upper limits are estimated for both fixed-radius and DR jet clustering methods. In all cases, the expected limit for the DR clustering is smaller than that for the fixed radius one, indicating that the analysis with DR jets has higher sensitivity to smaller coupling values. The improvement is more prominent for the fat jet $\text{p}_{\text{T}}$ threshold of 900 GeV, suggesting that the DR jet clustering is particularly effective for reconstructing signal events in the high fat jet $\text{p}_{\text{T}}$ regime, as indicated earlier in Section IV.
Figure 10: Comparison of $\text{FOM}_{\text{Punzi}}$ (left) and signal significance (right) between fixed-radius and dynamic radius clustering methods as a function of the fat jet $\text{p}_{\text{T}}$ threshold for the benchmark scenario defined in Eq. (7).
@@
Figure 11: Expected $2\sigma$ upper limit on the chromomagnetic coupling $C_{tRL}=C_{tLR}$ as a function of the top partner mass. The three panels correspond to three different fat jet $\text{p}_{\text{T}}$ thresholds. For these plots, we use $\Lambda=8$~TeV, and all other parameters are kept the same as in Eq. (7). Shaded regions indicate the parameter space excluded by our analysis, assuming no signal is observed.
@@ VI Summary and Outlook
In this work, we performed a thorough investigation of collider prospects of a vector-like top partner produced in association with a standard model top quark via a chromomagnetic coupling. The top partner is assumed to decay into the following channels: tZ, th, tg, or bW. To effectively suppress the huge QCD background, we focused on semileptonic final states, i.e., events with exactly one isolated lepton arising either from the top quark or from the heavy boson decay. Depending on the decay modes of the top partner, the final state could contain one or two fat jets along with a single lepton and b-tagged jet(s). We performed the analysis in the event category that contains a single fat jet ($\ell + 1$ fat jet + $\geqslant 1$ b-tagged jet) due to its comparatively better signal-to-background ratio.
We carried out a detailed collider analysis at the $\sqrt{s}=13$~TeV with an integrated luminosity of 300 fb-1, incorporating realistic pileup conditions. For signal events, the Madgraph5_aMC@NLO + Pythia8 + Delphes simulation chain has been employed; while for background estimation, appropriate simulated samples from CMS Open Data are utilized. To enhance signal discrimination, we utilized state-of-the-art jet substructure tools such as Soft Drop grooming, and $N$-subjettiness, along with several event shape observables like Fox-Wolfram moments, Sphericity, and Aplanarity. Multivariate discriminants based on boosted decision trees are developed for better signal-to-background separation. We then optimized a selection thresholds on these discriminants based on the Punzi figure-of-merit to improve signal sensitivity. The result is interpreted as the upper limit on the chromomagnetic coupling between the top quark and its heavy vector-like partner. Assuming the cut off scale of $\Lambda=8$ TeV, the upper limit on the coupling is found to lie within 0.2 to 1.5 for the top partner mass range of 1.4 to 3 TeV.